# Koopmon trajectories in nonadiabatic quantum-classical dynamics

### 21 Dec 2023 v1

@@
[…]
${}^{1}$epartment of Mathematics, University of Surrey, Guildford, United Kingdom
${}^{2}$ivision of Mathematical Sciences, Nanyang Technological University, Singapore
@@ 1.1 Trajectory-based schemes in molecular dynamics
In quantum molecular dynamics, several trajectory-based algorithms have been proposed over the years to overcome the computational challenges of fully quantum simulations. While many of these algorithms are traditionally based on the remarkable properties of Gaussian wavepackets [33, 46], a considerable effort has also been devoted to exploiting Madelung’s hydrodynamics [3, 20, 43, 67], which has the great advantage of restoring the concept of Lagrangian trajectory in quantum mechanics. Severe difficulties, however, challenge both approaches. For example, in the first case, the appearance of generally singular matrices often leads to uncontrolled approximations with the purpose of eliminating the degeneracy [60]. As a result [41], “the existent formalisms using classical frozen-width Gaussian motion do not conserve the total energy” and also struggle with momentum balance and conservation of probability [31].
@@
At the heart of the Born-Oppenheimer adiabatic approximation, this mixed quantum-classical (MQC) description represents an attractive avenue also in the more general context of nonadiabatic molecular dynamics. Yet, new problems emerge in extending the MQC method. For example, even if one discards the nuclear quantum potential, the need to retain correlation effects leads to dealing with the so-called quantum geometric tensor [19, 57], whose computational complexity is comparable to that of the quantum potential itself. For these reasons, MQC models continue to represent a subject of research in computational chemistry. Other areas in which the MQC description is attractive range from the theory of gravity [9] to spintronics [55], and solid-state physics [6, 38]. A possible realistic realization of an MQC system is represented by the molecular reaction dynamics occurring within a liquid solvent, for which the classical fluid description applies with great accuracy [14, 23].
@@ 1.2 Consistency of MQC theories and the Ehrenfest model
Here, $\widehat{H}(q,p)$ is a Hamiltonian (Hermitian) operator on the quantum Hilbert space $\mathscr{H}$ parameterized by the classical phase-space coordinates and $\hbar$ the reduced Planck constant. Also, $\widehat{\cal P}$ is a sufficiently smooth distribution taking values in the space $\operatorname{Her}(\mathscr{H})$ of Hermitian operators on $\mathscr{H}$. The classical Liouville density and the quantum density matrix are given as
@@ 1.3 Beyond the Ehrenfest model
The Koopman-von Neumann equation (KvN) of classical mechanics is written in terms of the canonical Poisson bracket as $i\hbar\partial_{t}\chi=\{i\hbar H,\chi\}$. Importantly, the latter identifies a linear unitary flow analogous to Schrödinger’s quantum mechanics. Due to the relation $f=|\chi|^{2}$, the KvN equation may be modified to include an arbitrary local phase factor, that is, $i\hbar\partial_{t}\chi=\{i\hbar H,\chi\}+\varphi\chi$ with $\varphi(q,p)$ being an arbitrary function. Since both quantum and classical dynamics can be formulated in terms of wavefunctions, the idea of considering their tensor-product space is rather intuitive and identifies the starting point of Sudarshan’s proposal. However, writing the dynamics for a MQC wavefunction $\Upsilon(q,p,x)$ on such a tensor-product space (here, $x$ is the quantum coordinate) is far from intuitive and we provide a summary of our derivation in the following section.
@@
We observe that both the vector field (6) and the effective Hamiltonian (7) are $\hbar$-corrections of the corresponding quantities appearing in the Ehrenfest model (2), that is $\langle{\mathbf{X}}_{\widehat{H}}\rangle$ and $\widehat{H}$, respectively. Recovering Born-Oppenheimer theory in the adiabatic approximation [7], this model was recently extended to consider MQC spin systems [21] and solute-solvent interaction in fluid solvents [23].
@@
The paper is organized as follows: In Section 2, we introduce the koopmon method. Starting from a brief review of the derivation and variational structure of the continuum model (5)-(7), we demonstrate how the particle method emerges form of regularization at the level of the underlying variational principle. Afterwards, in Section 3, we discuss the implementation of the koopmon method, which was used to produce the numerical experiments. In particular, we discuss its connection to the quantum bohmion method, and provide important details regarding the quantum solver and the output visualization. In Section 4, we present the results for the Tully models (I-III), which are well-known benchmark models in computational chemistry. The results for the Rabi models (ultrastrong and deep strong coupling) are presented in the following Section 5. In Section 6, we discuss the extension of both the koopmon and the bohmion method to higher dimensions. Finally, we present a brief conclusion and outlook in Section 7.
@@ 2.1 From Koopman wavefunctions to MQC dynamics
Once the classical phase $S({\mathbf{z}},t)$ is identified by the polar form $\chi=\sqrt{f}e^{iS/\hbar}$, applying the symmetry principle amounts to eliminating this phase from the action principle $\delta\int_{t_{1}}^{t_{2}}\!\int\!f(\partial_{t}S-{\langle\psi,i\hbar\partial_{t}\psi-\{i\hbar\widehat{H},\psi\}+(p\partial_{p}\widehat{H}-\widehat{H})\psi+\{S,{\widehat{H}}\}\psi\rangle})\,{\rm d}^{2}z=0$. Benefiting from the properties of the Berry connection $\langle\psi|-i\hbar\nabla\psi\rangle$ [8], this step requires introducing diffeomorphic Lagrangian paths $\boldsymbol{\eta}_{t}\in\operatorname{Diff}(T^{*}Q)$ on phase space, which are generated by the vector field $\boldsymbol{\cal X}$ via the relation $\dot{\boldsymbol{\eta}_{t}}=\boldsymbol{\cal X}\circ\boldsymbol{\eta}_{t}$. Here, $\operatorname{Diff}(T^{*}Q)$ denotes the group of diffeomorphism on the classical phase space. In particular, one combines integration by parts with the continuity equation $\partial_{t}f+\operatorname{div}(f\boldsymbol{\cal X})=0$ to write $\delta\int_{t_{1}}^{t_{2}}\!\int\!f\partial_{t}S\,{\rm d}^{2}z{\rm d}t=-\delta\int_{t_{1}}^{t_{2}}\!\int\!f\nabla S\cdot\boldsymbol{\cal X}\,{\rm d}^{2}z{\rm d}t$, where $f$ and $\boldsymbol{\cal X}$ are now varied as in Euler-Poincaré theory [35], that is $\delta f=-\operatorname{div}(f\boldsymbol{\cal Y})$ and $\delta\boldsymbol{\cal X}=\partial_{t}\boldsymbol{\cal Y}+\boldsymbol{\cal X}\cdot\nabla\boldsymbol{\cal Y}-\boldsymbol{\cal Y}\cdot\nabla\boldsymbol{\cal X}$, where $\boldsymbol{\cal Y}=\delta\boldsymbol{\eta}_{t}\circ\boldsymbol{\eta}_{t}^{-1}$ is arbitrary. Finally, one last convenient step consists in writing $\psi(t)=(U_{t}\psi_{0})\circ\boldsymbol{\eta}_{t}^{-1}$, without loss of generality [22]. Here, $U_{t}=U_{t}({\mathbf{z}})$ is a unitary operator on the quantum Hilbert space that is parameterized by phase-space coordinates. This last step simply expresses the quantum unitary dynamics in the frame of Lagrangian classical paths. In this way, the action principle is entirely expressed in terms of $f$ and $\widehat{\cal P}=\Upsilon\Upsilon^{\dagger}=f\psi\psi^{\dagger}$ as
@@
$\delta\!\int_{t_{1}}^{t_{2}}\!\!\left(\int\!\big{(}f\boldsymbol{\cal A}\cdot\boldsymbol{\cal X}+\langle\widehat{\cal P},i\hbar\hat{\xi}\rangle\big{)}{\rm d}^{2}z-h(f,\widehat{\cal P})\!\right)\!{\rm d}t=0,\qquad\quad h(f,\widehat{\cal P})=\!\int\!\big{\langle}f{\widehat{H}}+i\hbar\{\widehat{\cal P},\widehat{H}\}\big{\rangle}\,{\rm d}^{2}z,$ (8)
@@
which is indeed phase-independent. Here, ${\boldsymbol{\cal A}=(p,0)}$ is the coordinate representation of the one-form ${\cal A}={\boldsymbol{\cal A}\cdot{\rm d}{\mathbf{z}}}=p{\rm d}q$. Also, $\hat{\xi}=(\dot{U}_{t}U_{t}^{\dagger})\circ\boldsymbol{\eta}_{t}^{-1}$ is skew-Hermitian, so that $\delta\hat{\xi}=\partial_{t}\hat{\Lambda}+[\hat{\Lambda},\hat{\xi}]+\boldsymbol{\cal X}\cdot\nabla\hat{\Lambda}-\boldsymbol{\cal Y}\cdot\nabla\hat{\xi}$ and $\delta\widehat{\cal P}=[\hat{\Lambda},\widehat{\cal P}]-\operatorname{div}(\widehat{\cal P}\boldsymbol{\cal Y})$, where $\hat{\Lambda}=(\delta{U}_{t}U_{t}^{\dagger})\circ\boldsymbol{\eta}_{t}^{-1}$ is skew-Hermitian and arbitrary. These variations arise by standard Euler-Poincaré reduction from Lagrangian to Eulerian variables [35]. Indeed, Lagrangian trajectories play a crucial role in the variational problem associated to (8). In particular, if $\boldsymbol{\eta}({\mathbf{z}}_{0},t)$ is the diffeomorphic Lagrangian path on phase space and $U({\mathbf{z}},t)$ is a unitary operator, the Eulerian quantities evolve as follows:
@@
where $\boldsymbol{\eta}_{t*}$ denotes the push-forward by the time-dependent map $\boldsymbol{\eta}_{t}$. As shown in [24, 62], the variational principle (8), with the expressions above for the Euler-Poincaré variations $\delta f$, $\delta\widehat{\cal P}$, $\delta\boldsymbol{\cal X}$, and $\delta\hat{\xi}$, return the model equation (5) along with its trace, that is $\partial_{t}{f}+\operatorname{div}\!\big{(}f\boldsymbol{\cal X}\big{)}=0$, and the accompanying relations (6)-(7). Notice that the total energy is identified by the Hamiltonian functional $h(f,\widehat{\cal P})$ in (8), which is given by the sum of usual average $\int\!f\langle\widehat{H}\rangle{\rm d}^{2}z$ with an extra term rendering the quantum backreaction, that is the dynamical features specifically associated to the statistical correlations between the quantum and the classical systems. Here, these correlations appear through the inhomogeneities encoded in the integrand term $\langle i\hbar\{\widehat{\cal P},\widehat{H}\}\rangle={f^{-1}\operatorname{Tr}({\mathbf{X}}_{\widehat{H}}\cdot[i\hbar\widehat{\cal P},\nabla\widehat{\cal P}])}/2$. As noted in [23], an analogous term also appears in the theory of spin-orbit coupling and this analogy hinges on the prominent role of the Mead connection $i\hbar f^{-2}[\widehat{\cal P},\nabla\widehat{\cal P}]$ [49]. This and related aspects will be developed elsewhere.
@@ 2.2 Regularization and the particle scheme
Given the challenging nature of the nonlinear PDEs (5)-(7), here we approximate their associated dynamics by devising a particle method that is based on a variational regularization. The treatment follows an analogous construction developed within our team in the context of quantum hydrodynamics [19, 36], and later applied to Born-Oppenheimer dynamics [59].
@@
[…]
with the same Euler-Poincaré variations as above. Note that regularized functions are denoted by a bar. At this point, taking the time derivative of the second in (9) yields
@@
Moreover, taking variations of (10) yields
@@
[…]
@@
Notice that discarding the last term in the expression above recovers the multi-trajectory Ehrenfest system (4). Indeed, both models share the same Hamiltonian structure that is identified by the direct-sum Poisson bracket
@@
Also, equations (14)-(15) inherit all the MQC consistency criteria listed in Section 1.2. Despite these analogies, we observe that the koopmon scheme differs substantially from (4) in that each particle is coupled to all others at all points in time. This coupling among trajectories reflects the quantum-classical correlation effects and indeed vanishes when $\widehat{H}$ is purely classical (that is $\widehat{H}=H\boldsymbol{1}$) or purely quantum (that is $\nabla\widehat{H}=0$).

### 17 Jun 2024 v2

@@
Werner Bauer1, Paul Bergold1111Corresponding author (p.bergold@surrey.ac.uk) , François Gay-Balmaz2, Cesare Tronci1
1Department of Mathematics, University of Surrey, Guildford, United Kingdom
2Division of Mathematical Sciences, Nanyang Technological University, Singapore
@@ 1.1 Trajectory-based schemes in molecular dynamics
In quantum molecular dynamics, several trajectory-based algorithms have been proposed over the years to overcome the computational challenges of fully quantum simulations. While many of these algorithms are traditionally based on the remarkable properties of Gaussian wavepackets [35, 49], a considerable effort has also been devoted to exploiting Madelung’s hydrodynamics [3, 21, 45, 77], which has the great advantage of restoring the concept of Lagrangian trajectory in quantum mechanics. Severe difficulties, however, challenge both approaches. For example, in the first case, the appearance of generally singular matrices often leads to uncontrolled approximations with the purpose of eliminating the degeneracy [68]. As a result [43], “the existent formalisms using classical frozen-width Gaussian motion do not conserve the total energy” and also struggle with momentum balance and conservation of probability [32]. Indeed, these limitations generally affect multiconfiguration methods and other approaches based on basis-set expansions; see [43] for further details.
@@
Arising naturally within the Born–Oppenheimer adiabatic approximation, this mixed quantum -classical (MQC) description represents an attractive avenue also in the more general context of nonadiabatic molecular dynamics. Yet, new problems emerge in extending the MQC method. For example, even if one discards the nuclear quantum potential, the need to retain correlation effects leads to dealing with the so-called quantum geometric tensor [20, 65], whose computational complexity is comparable to that of the quantum potential itself. For these reasons, MQC models continue to represent a subject of research in computational chemistry. Other areas in which the MQC description is attractive range from the theory of gravity [9] to spintronics [63], and solid-state physics [6, 40]. A possible realistic realization of an MQC system is represented by the molecular reaction dynamics occurring within a liquid solvent, for which the classical fluid description applies with great accuracy [15, 24].
@@ 1.2 Consistency of MQC theories and the Ehrenfest model
Here, $\widehat{H}(q,p)$ is a Hamiltonian (Hermitian) operator on the quantum Hilbert space $\mathscr{H}$ parameterized by the classical phase-space coordinates and $\hbar$ denotes the reduced Planck constant. Also, $\widehat{\cal P}$ is a sufficiently smooth distribution taking values in the space $\operatorname{Her}(\mathscr{H})$ of Hermitian operators on $\mathscr{H}$. The classical Liouville density and the quantum density matrix are given as
@@ 1.3 Beyond the Ehrenfest model
The Koopman-von Neumann equation (KvN) of classical mechanics is written in terms of the canonical Poisson bracket as $i\hbar\partial_{t}\chi=\{i\hbar H,\chi\}$. The latter means that the classical evolution of Koopman wavefunctions is given as $\chi(t)=\chi_{0}\circ{\boldsymbol{\eta}}_{t}^{-1}$, that is the composition of the initial condition $\chi_{0}$ by the inverse of the Lagrangian map ${\boldsymbol{\eta}}_{t}$ generated by the vector field ${\mathbf{X}}_{H}$. We remark that the composition map is also known as the Koopman operator and has recently attracted a great deal of attention in dynamical systems and data science [46, 50, 56].
Importantly, the KvN equation identifies a linear unitary flow analogous to Schrödinger’s quantum mechanics. Due to the relation $f=|\chi|^{2}$, the KvN equation may be modified to include an arbitrary local phase factor, that is, $i\hbar\partial_{t}\chi=\{i\hbar H,\chi\}+\varphi\chi$ with $\varphi(q,p)$ being an arbitrary function. Since both quantum and classical dynamics can be formulated in terms of wavefunctions, the idea of considering their tensor-product space is rather intuitive and identifies the starting point of Sudarshan’s proposal. However, writing the dynamics for a MQC wavefunction $\Upsilon(q,p,x)$ on such a tensor-product space (here, $x$ is the quantum coordinate) is far from intuitive and we provide a summary of our derivation in the following section.
@@
We observe that both the vector field (6) and the effective Hamiltonian (7) are $\hbar$-corrections of the corresponding quantities appearing in the Ehrenfest model (2), that is $\langle{\mathbf{X}}_{\widehat{H}}\rangle$ and $\widehat{H}$, respectively. Recovering Born–Oppenheimer molecular dynamics in the adiabatic approximation [7], this model was recently extended to consider MQC spin systems [22] and solute-solvent interaction in fluid solvents [24]. The corresponding Heisenberg picture is also available in [54].
@@
The paper is organized as follows: In Section 2, we introduce the koopmon method. Starting from a brief review of the derivation and variational structure of the continuum model (5)-(7), we demonstrate how the particle method emerges form of regularization at the level of the underlying variational principle. Afterwards, in Section 3, we discuss the implementation of the koopmon method, which was used to produce the numerical experiments. In particular, we discuss its connection to the quantum bohmion method, and provide important details regarding the quantum solver and the output visualization. In Section 4, we present the results for the Tully models (I-III), which are well-known benchmark models in computational chemistry. The results for the Rabi models (ultrastrong and deep strong coupling) are presented in the following Section 5. In Section 6, we discuss the extension of both the koopmon and the bohmion method to higher dimensions. Finally, we present a brief conclusion and outlook in Section 7, while Appendix A expands on certain concepts commonly used in the quantum chemistry of nonadiabatic processes.
@@ 2.1 From Koopman wavefunctions to MQC dynamics
Once the classical phase $S({\mathbf{z}},t)$ is identified by the polar form $\chi=\sqrt{f}e^{iS/\hbar}$, applying the symmetry principle amounts to eliminating this phase from the action principle $\delta\int_{t_{1}}^{t_{2}}\!\int\!f(\partial_{t}S-{\langle\psi,i\hbar\partial_{t}\psi-\{i\hbar\widehat{H},\psi\}+(p\partial_{p}\widehat{H}-\widehat{H})\psi+\{S,{\widehat{H}}\}\psi\rangle})\,{\rm d}^{2}z=0$. Benefiting from the properties of the Berry connection $\langle\psi|-i\hbar\nabla\psi\rangle$ [8], this step requires introducing diffeomorphic Lagrangian paths $\boldsymbol{\eta}_{t}\in\operatorname{Diff}(T^{*}Q)$ on phase space, which are generated by the vector field $\boldsymbol{\cal X}$ via the relation $\dot{\boldsymbol{\eta}_{t}}=\boldsymbol{\cal X}\circ\boldsymbol{\eta}_{t}$. Here, $\operatorname{Diff}(T^{*}Q)$ denotes the group of diffeomorphism on the classical phase space. In particular, one combines integration by parts with the continuity equation $\partial_{t}f+\operatorname{div}(f\boldsymbol{\cal X})=0$ to write $\delta\int_{t_{1}}^{t_{2}}\!\int\!f\partial_{t}S\,{\rm d}^{2}z{\rm d}t=-\delta\int_{t_{1}}^{t_{2}}\!\int\!f\nabla S\cdot\boldsymbol{\cal X}\,{\rm d}^{2}z{\rm d}t$, where $f$ and $\boldsymbol{\cal X}$ are now varied as in Euler–Poincaré theory [37], that is $\delta f=-\operatorname{div}(f\boldsymbol{\cal Y})$ and $\delta\boldsymbol{\cal X}=\partial_{t}\boldsymbol{\cal Y}+\boldsymbol{\cal X}\cdot\nabla\boldsymbol{\cal Y}-\boldsymbol{\cal Y}\cdot\nabla\boldsymbol{\cal X}$, where $\boldsymbol{\cal Y}=\delta\boldsymbol{\eta}_{t}\circ\boldsymbol{\eta}_{t}^{-1}$ is arbitrary. Finally, one last convenient step consists in writing $\psi(t)=(U_{t}\psi_{0})\circ\boldsymbol{\eta}_{t}^{-1}$, without loss of generality [23]. Here, $U_{t}=U_{t}({\mathbf{z}})$ is a unitary operator on the quantum Hilbert space that is parameterized by phase-space coordinates. This last step simply expresses the quantum unitary dynamics in the frame of Lagrangian classical paths. In this way, the action principle is entirely expressed in terms of $f$ and $\widehat{\cal P}=\Upsilon\Upsilon^{\dagger}=f\psi\psi^{\dagger}$ as
@@
$\delta\!\int_{t_{1}}^{t_{2}}\!\!\left(\int\!\big{(}f\boldsymbol{\cal A}\cdot\boldsymbol{\cal X}+\langle\widehat{\cal P},i\hbar\hat{\xi}\rangle\big{)}{\rm d}^{2}z-h(f,\widehat{\cal P})\!\right)\!{\rm d}t=0,\quad\,\,\,\,\,h(f,\widehat{\cal P})=\!\int\!\big{\langle}f{\widehat{H}}+i\hbar\{\widehat{\cal P},\widehat{H}\}\big{\rangle}\,{\rm d}^{2}z,$ (8)
@@
which is indeed phase-independent. Here, ${\boldsymbol{\cal A}=(p,0)}$ is the coordinate representation of the canonical one-form ${\cal A}={\boldsymbol{\cal A}\cdot{\rm d}{\mathbf{z}}}=p{\rm d}q$. Also, $\hat{\xi}=(\dot{U}_{t}U_{t}^{\dagger})\circ\boldsymbol{\eta}_{t}^{-1}$ is skew-Hermitian, so that $\delta\hat{\xi}=\partial_{t}\hat{\Lambda}+[\hat{\Lambda},\hat{\xi}]+\boldsymbol{\cal X}\cdot\nabla\hat{\Lambda}-\boldsymbol{\cal Y}\cdot\nabla\hat{\xi}$ and $\delta\widehat{\cal P}=[\hat{\Lambda},\widehat{\cal P}]-\operatorname{div}(\widehat{\cal P}\boldsymbol{\cal Y})$, where $\hat{\Lambda}=(\delta{U}_{t}U_{t}^{\dagger})\circ\boldsymbol{\eta}_{t}^{-1}$ is skew-Hermitian and arbitrary. These variations arise by standard Euler–Poincaré reduction from Lagrangian to Eulerian variables [37]. Indeed, Lagrangian trajectories play a crucial role in the variational problem associated to (8). In particular, if $\boldsymbol{\eta}({\mathbf{z}}_{0},t)$ is the diffeomorphic Lagrangian path on phase space and $U({\mathbf{z}},t)$ is a unitary operator, the Eulerian quantities evolve as follows:
@@
where $\boldsymbol{\eta}_{t*}$ denotes the push-forward by the time-dependent map $\boldsymbol{\eta}_{t}$. As shown in [25, 72], the variational principle (8), with the expressions above for the Euler–Poincaré variations $\delta f$, $\delta\widehat{\cal P}$, $\delta\boldsymbol{\cal X}$, and $\delta\hat{\xi}$, return the model equation (5) along with its trace, that is $\partial_{t}{f}+\operatorname{div}\!\big{(}f\boldsymbol{\cal X}\big{)}=0$, and the accompanying relations (6)-(7). Notice that the total energy is identified by the Hamiltonian functional $h(f,\widehat{\cal P})$ in (8), which is given by the sum of usual average $\int\!f\langle\widehat{H}\rangle{\rm d}^{2}z$ with an extra term rendering the quantum backreaction, that is the dynamical features specifically associated to the statistical correlations between the quantum and the classical systems. Here, these correlations appear through the inhomogeneities encoded in the integrand term $\langle i\hbar\{\widehat{\cal P},\widehat{H}\}\rangle={f^{-1}\operatorname{Tr}({\mathbf{X}}_{\widehat{H}}\cdot[i\hbar\widehat{\cal P},\nabla\widehat{\cal P}])}/2$ in (8). As noted in [24, 54], an analogous term also appears in the theory of spin-orbit coupling and this analogy hinges on the prominent role of the Mead connection $i\hbar f^{-2}[\widehat{\cal P},\nabla\widehat{\cal P}]$ [55]. This and related aspects will be developed elsewhere.
@@ 2.2 Regularization and the particle scheme
Given the challenging nature of the nonlinear PDEs (5)-(7), here we approximate their associated dynamics by devising a particle method that is based on a variational regularization. The treatment follows an analogous construction developed within our team in the context of quantum hydrodynamics [20, 38], and later applied to Born–Oppenheimer dynamics [67].
@@
$\delta\!\int_{t_{1}}^{t_{2}}\!\!\left(\int\!\big{(}f\boldsymbol{\cal A}\cdot\boldsymbol{\cal X}+\langle\widehat{\cal P},i\hbar\hat{\xi}-\widehat{H}\rangle\big{)}{\rm d}^{2}z-{\cal B}_{\textit{reg}}\right)\!{\rm d}t=0,\quad\,\,\,\,{\cal B}_{\textit{reg}}=\!\int\!\bar{f}^{-1}\langle\bar{\cal P},i\hbar\{\bar{\cal P},\widehat{H}\}\big{\rangle}\,{\rm d}^{2}z,$ (10)
with the same Euler–Poincaré variations as above. Note that regularized functions are denoted by a bar. At this point, taking the time derivative of the second in (9) yields
@@
Moreover, taking variations in (10) yields
@@
Here, we have used the standard notation for functional derivatives as well as the chain-rule relations ${\delta{\cal B}_{\textit{reg}}}/{\delta f}=\overline{{\delta{\cal B}(\bar{f},\bar{\cal P})}/{\delta\bar{f}}}$ and ${\delta{\cal B}_{\textit{reg}}}/{\delta\widehat{\cal P}}=\overline{{\delta{\cal B}(\bar{f},\bar{\cal P})}/{\delta\bar{\cal P}}}$, resulting from ${\cal B}_{\textit{reg}}(f,\widehat{\cal P})={\cal B}(\bar{f},\bar{\cal P})$. As a result, the singular ansatz (3) is now an exact solution of the equation ${\partial_{t}}f+\operatorname{div}(f\boldsymbol{{\cal X}})=0$, so that $\dot{{\boldsymbol{\zeta}}}_{a}=\boldsymbol{{\cal X}}({\boldsymbol{\zeta}}_{a})$ and one can proceed similarly to the discussion in Section 1.2 upon writing the equation $\partial_{t}\,\widehat{\!\mathscr{P}}+\boldsymbol{{\cal X}}\cdot\nabla\,\widehat{\!\mathscr{P}}=[\hat{\xi},\,\widehat{\!\mathscr{P}}]$ for the quantity $\,\widehat{\!\mathscr{P}}=\widehat{\cal P}/f$. This procedure leads again to a set of closed equations for $({\boldsymbol{\zeta}}_{a}(t),\hat{\varrho}_{a}(t))$, where $\hat{\varrho}_{a}=\,\widehat{\!\mathscr{P}}({\boldsymbol{\zeta}}_{a})$, thereby extending the multi-trajectory Ehrenfest scheme (4). Given the complexity of the regularized continuum model (11)-(12), however, here we follow a different, yet equivalent method.
@@
Notice that discarding the last term in the expression above recovers the multi-trajectory Ehrenfest system (4). Indeed, both the latter and the koopmon model share the same Hamiltonian structure that is identified by the direct-sum Poisson bracket
@@
Also, equations (14)-(15) inherit all the MQC consistency criteria listed in Section 1.2. Despite these analogies, we observe that the koopmon scheme differs substantially from (4) in that each particle is coupled to all others at all points in time. This coupling among trajectories reflects the quantum-classical correlation effects and indeed vanishes when $\widehat{H}$ is purely classical (that is $\widehat{H}=H\boldsymbol{1}$) or purely quantum (that is $\nabla\widehat{H}=0$). Also, we observe that this coupling does not take place via additional force terms in a standard set of Newtonian equations. Instead, the coupling occurs via an integral involving both the position and momentum of all particles, thereby affecting the evolution of both classical variables.
As previously illustrated in the plan of the paper, so far we have reviewed the main differences between Ehrenfest dynamics and our new MQC model in (5)-(7). In particular, after summarizing the formulation of the underlying continuum PDEs, we have proceeded by presenting the corresponding particle closures. The reminder of this work presents the results obtained by the implementation of the koopmon scheme along, with a comparison with analogous trajectory methods including the multi-trajectory Ehrenfest method.